2x + 3y=13 and x - y=9
Answer by rfer(16322)

x-y=9, x=y+9
2x+3y=13
2(y+9)+3y=13
2y+18+3y=13
5y=-5
y=-1
-----------------
x-(-1)1=9
x8
--------------
2(8)+3(-1)=13